solve for substition x=-1+y
23x-7y=-7
find the solution of this system of equations. find what x= and what y=

Answer by Maths68(1474)

x=-1+y...............(1)
23x-7y=-7............(2)
Put the value of x from (1) to (2)
23(-1+y)-7y=-7
-23+23y-7y=-7
23y-7y=-7+23
16y=16
16y/16=16/16
y=1
Put the value of y in (1)
x=-1+1
x=0
x,y=0,1